OpenStudy (anonymous):

can someone solve this for me and show their work step by step, please ? derivative of radical{(x-1)/(x+1)}

OpenStudy (unklerhaukus):

\[\sqrt\frac{x-1}{x+1}\]To rationalise the denominator Multiply by a special form of one \[=\sqrt\frac{x-1}{x+1}\times1\\ =\frac{\sqrt{x-1}}{\sqrt{x+1}}\times\frac{\sqrt{x+1}}{\sqrt{x+1}}\]

OpenStudy (anonymous):

i' sorry, i forgot to type that I was looking for the derivative

OpenStudy (unklerhaukus):

oh right

OpenStudy (unklerhaukus):

you could use the quotient rule (that's always fun) \[\left(\frac{u}{v}\right)'=\frac{u'v-v'u}{v^2}\]

OpenStudy (unklerhaukus):

There is no reason to rationalise the denominator now \[u=\sqrt{x-1}\implies u'=\]\[v=\sqrt{x+1}\implies v'=\\ \qquad\qquad\quad\implies v^2=\]

OpenStudy (anonymous):

Thanks!

OpenStudy (unklerhaukus):

and it might be easier if you express the square roots as indices ^(1/2)